A train travelling at 12 m/s accelerates to 20 m/s in 5 seconds.

What will its speed (final velocity) be after another 3 seconds if it continues to accelerate at the same rate correct to 1 decimal place?


1
Expert's answer
2020-08-19T12:56:58-0400

The acceleration of a train

"a=\frac{\Delta v}{\Delta t}=\frac{(20-12)\:\rm m/s}{5\:\rm s}=1.6\:\rm m/s^2"

The final velocity

"v_f=v_i+at\\
=20\:\rm m/s+1.6\: m/s^2\times 3\: s=24.8\: m/s"